jigglypuff314 (jigglypuff314):

Check my answers? :3 f(t) = 6 + cos(t/10) + 3sin(7t/40) is used to model the velocity of the plane, in miles per minute, for 0 ≤ t ≤ 40 what is the total distance traveled by the plane over the time interval 0 ≤ t ≤ 40? (I got 236.651 miles ???) If at time t=5 the test plane is at mile marker 15 along the straight test track, determine the position of the plane at time t=25. (I got 161.659 miles ???)

OpenStudy (anonymous):

JIGGLY! LONG TIME NO SEE

ganeshie8 (ganeshie8):

you can do that cuz the given velocity function is never negative

jigglypuff314 (jigglypuff314):

yeah, even then all you would have to do is put abs value in there somewhere XD :P thanks :)

jigglypuff314 (jigglypuff314):

how about for the other part? :3

ganeshie8 (ganeshie8):

the other part requires taking indefinite integral, solving for the constant and finding d(25) uhmm

ganeshie8 (ganeshie8):

let me check quick...

jigglypuff314 (jigglypuff314):

6t + 10sin(t/10) - (120/7)cos(7t/40) + C plugged in t=5 found C= 1.194 ? plugged in t=25 -> 161.658 ?

ganeshie8 (ganeshie8):

\(\large d(t) = 6t + 10\sin(\frac{t}{10}) - \frac{120}{7}\cos(\frac{7t}{40})+c\)

ganeshie8 (ganeshie8):

check once.. .

jigglypuff314 (jigglypuff314):

oh yeah I see what I did wrong, I made it equal 25 instead of 15 :P whoops

ganeshie8 (ganeshie8):

ohk :)

jigglypuff314 (jigglypuff314):

so plugged in 25 -> 152.854 ??? :3

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

jigglypuff314 (jigglypuff314):

great! :D Thanks!!! :)

ganeshie8 (ganeshie8):

np :)